1.How many grams of potassium chloride, KCl, are produced if 50.0g of potassium chlorate, KClO3, decompose?
2KClO3 → 2KCl + 3O2
2.How many grams of hydrogen, H2, are necessary to react completely with
20.0 g of nitrogen, N2?
N2 + 3H2 → 2NH3
3.How many grams of AgCl, silver chloride, are produced from 65.0 g of AgNO3, silver nitrate?
2AgNO3 + BaCl2 → 2AgCl + Ba(NO3) 2
1
Expert's answer
2020-12-06T13:29:44-0500

1. M(KClO3) = 122.55 g/mol

n(KClO3) "= \frac{50.0}{122.55} = 0.408 \;mol"

n(KCl) = n(KClO3) = 0.408 mol

M(KCl) = 74.55 g/mol

m(KCl) "= 0.408 \times 74.55 = 30.41 \;g"

2. M(N2) = 28.01 g/mol

n(N2) "= \frac{20.0}{28.01} = 0.714 \;mol"

n(H2) = 3n(N2) "= 3 \times 0.714 = 2.142 \;mol"

M(H2) = 2 g/mol

m(H2) "= 2 \times 2.142 = 4.284 \;g"

3. M(AgNO3) = 169.87 g/mol

n(AgNO3) "= \frac{65.0}{169.87} = 0.382 \;mol"

n(AgCl) = n(AgNO3) = 0.382 mol

M(AgCl) = 143.32 g/mol

m(AgCl) = "0.382 \times 143.32 = 54.75 \;g"